# Analytic Models for the Capacity Distribution in MDG-impaired Optical SDM Transmission

### v2

@@ V Impact of frequency diversity
Figure 9: Total capacity deviation per mode count for ${\sigma_{mdg}=5~{}\mathrm{dB}}$ from the analytical equations.
So far, this paper considered a narrowband channel with a flat frequency response. However, as presented in [15], modal dispersion induces frequency-dependent mode gains. In this case, the total capacity in a given bandwidth can be calculated by dividing the spectrum into $N$ narrowband frequency bins and adding the corresponding capacities [16]. In this model, the channel response of a given frequency bin is independent of that in adjacent bins [15]. Considering the distributions presented in Section III, the averaged capacity variance of the $i^{th}$ mode, considering $N$ independent frequency bins is given by
@@ VI Case Study
The $\lambda_{\mathrm{dB}}$ distribution coefficients can be obtained from [14, Table 1] or derived analytically, as detailed in appendix A. They are given by
@@
Figure 10: Total Gaussian PDF for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{mdg}=5~{}\mathrm{dB}}$. For the frequency diverse case, we consider 2 uncorrelated frequency bins. The estimated outage capacities consider an outage probability of ${p_{\mathrm{out}}=1\%}$.
@@ Appendix A Distribution of the ensemble of $\lambda_{\mathrm{dB}}$ eigenvalues
As described in [22], the logarithm power gain $\lambda_{\mathrm{dB}}$ follows the fixed-trace GUE. The joint PDF of the GUE spectral distribution is given by [27, Eq. (3.3.7)]
@@
The ensemble spectral distribution, related to any eigenvalue $\lambda$ in set $\lambda_{i},i\in[1,D]$, is obtained from the $k$-point correlation function in the case of $k=1$, defined in [29, III] and [27, Eq. (6.1.2)].
The fixed-trace solution of the ensemble spectral distribution normalized to zero mean is developed in detail in [22, From (S.3) to (S.15)]. Defining an auxiliary variable $t$ where its $n^{th}$ order polynomial is expressed in terms of $\lambda$ as
@@
As pointed out by [20, I], at the limit $D\rightarrow\infty$ the distribution given in (38) converges to
@@
Appendix B Equation for the mean $\mu_{\lambda_{\mathrm{dB}}}$ of the linearized eigenvalues
Given $\lambda_{\mathrm{dB}}=10\cdot\mathrm{log}_{10}\left(\lambda\right)$, the linearized eigenvalue distribution is related to the $\lambda_{\mathrm{dB}}$ distribution by
@@ Appendix C Approximation of $\mu_{\lambda_{\mathrm{dB},i}}$ from the GUE derivative
Setting $g(\widehat{\lambda}_{\mathrm{dB}})=\sum_{k=0}^{D-1}\beta_{\lambda_{\mathrm{dB}},D,k}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2k}}{\sigma_{\mathrm{mdg}}^{2k}}$ and $\delta_{0}=\frac{D+1}{2\sigma_{mdg}^{2}}$,

### v3

@@ V Impact of frequency diversity
Figure 10: Total capacity deviation per mode count for ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$ from the analytical equations.
While this paper has so far only discussed a narrowband channel, as presented in [24], modal dispersion induces frequency-dependent mode gains. In this case, the total capacity in a given bandwidth can be calculated by dividing the spectrum into $N$ narrowband frequency bins with uncorrelated modal gains, i.e. $N$ subchannels, and adding the corresponding capacities [25]. In this model, the channel response of a given frequency bin is independent of that in adjacent bins [24]. As the coherence bandwidth of the modal gains is in the order of the inverse standard deviation of the group delay $\sigma_{\text{GD}}$ [24], $N$ is linearly proportional to $\sigma_{\text{GD}}$. Indeed, $N$ can be approximated by the given bandwidth multiplied by $\sigma_{\text{GD}}$ [24, 33] Considering the distributions presented in Section III, the averaged capacity variance of the $i^{th}$ mode, considering $N$ independent frequency bins is given by
@@ VI Case Study
The $\lambda_{\mathrm{dB}}$ distribution coefficients can be obtained from [23, Table 1] or derived analytically, as detailed in appendix A. They are given by
@@
We validate the frequency diverse case against simulated values in Fig. 11. For each Monte-Carlo trial, we simulate 2 uncorrelated channel matrices $\mathbf{H}$, and average the capacity between both subchannels.
Figure 11: Total Gaussian PDF for ${D=6}$, ${\mathrm{SNR}=10~{}\mathrm{dB}}$, and ${\sigma_{\mathrm{mdg}}=5~{}\mathrm{dB}}$. For the frequency diverse case, we consider 2 uncorrelated frequency bins, where we show the capacity average over the subchannels. The estimated outage capacities consider an outage probability of ${p_{\mathrm{out}}=1\%}$.
@@ Appendix A Distribution of the ensemble of $\lambda_{\mathrm{dB}}$ eigenvalues
As described in [31], the logarithm power gain $\lambda_{\mathrm{dB}}$ follows the fixed-trace GUE. The joint PDF of the GUE spectral distribution is given by [36, Eq. (3.3.7)]
@@
The ensemble spectral distribution, related to any eigenvalue $\lambda$ in set $\lambda_{i},i\in[1,D]$, is obtained from the $k$-point correlation function in the case of $k=1$, defined in [38, III] and [36, Eq. (6.1.2)].
The fixed-trace solution of the ensemble spectral distribution normalized to zero mean is developed in detail in [31, From (S.3) to (S.15)]. Defining an auxiliary variable $t$ where its $n^{th}$ order polynomial is expressed in terms of $\lambda$ as
@@
As pointed out by [32, I], at the limit $D\rightarrow\infty$ the distribution given in (38) converges to
@@
Appendix B Equation for the mean $\mu_{\lambda_{\mathrm{dB}}}$ of the logarithmic scale eigenvalues $\lambda_{\mathrm{dB}}$
Given $\lambda_{\mathrm{dB}}=10\cdot\mathrm{log}_{10}\left(\lambda\right)$, the linear scale eigenvalue distribution is related to the $\lambda_{\mathrm{dB}}$ distribution by
@@ Appendix C Approximation of $\mu_{\lambda_{\mathrm{dB},i}}$ from the GUE derivative
Setting $g(\widehat{\lambda}_{\mathrm{dB}})=\sum_{k=0}^{D-1}\beta_{\lambda_{\mathrm{dB}},D,k}\frac{\widehat{\lambda}_{\mathrm{dB}}^{2k}}{\sigma_{\mathrm{mdg}}^{2k}}$ and $\delta_{0}=\frac{D+1}{2\sigma_{\mathrm{mdg}}^{2}}$,